# AnywhereDoor: Multi-Target Backdoor Attacks on Object Detection

### v1

@@ 4.1 Objective Disentanglement
The essence of backdoor attacks on object detection models lies in exploiting DNNs’ excessive learning ability to link triggers with malicious outputs [33]. Existing backdoor attacks on object detection [2, 8, 41] rely on fixed trigger patterns, limiting the attack to a single predefined behavior. This limitation arises due to the exponential growth in pattern-prediction associations when introducing varied triggers and behaviors, which exceeds the network’s learning capacity, drastically reducing ASR by making it harder to differentiate between the many pattern-prediction pairs.
Our proposed strategy, objective disentanglement, decouples the attack into removal and generation components, which can constitute the five attack scenarios. They are represented by two sub-vectors that make up the intent embeddings, as illustrated in Fig. 4, each being a zero-filled, one-filled or one-hot vector with a length matching the number of classes. The trigger generator is similarly divided into two sub-networks, each mapping one sub-vector to a trigger.
Formally, let $\mathbf{e}=[\mathbf{e}_{r},\mathbf{e}_{g}]$ be the intent embedding, where $\mathbf{e}_{r}$ and $\mathbf{e}_{g}$ are the sub-vectors for removal and generation, respectively. The trigger generator $G_{\boldsymbol{\phi}}$ is divided into two sub-networks $G_{\boldsymbol{\phi}_{r}}$ and $G_{\boldsymbol{\phi}_{g}}$, such that: $\mathbf{t}_{r}=G_{\boldsymbol{\phi}_{r}}(\mathbf{e}_{r})$ and $\mathbf{t}_{g}=G_{\boldsymbol{\phi}_{g}}(\mathbf{e}_{g})$. The final trigger $\mathbf{t}$ is obtained by combining their outputs: $\mathbf{t}=\mathbf{t}_{r}+\mathbf{t}_{g}$. This trigger unit is then applied to produce the poisoned image utilizing the trigger mosaicking technique as discussed by Sec. 4.2.
Figure 4: Objective disentanglement illustration. The intent embeddings are splitted into two parts that represent the attacker’s intent collectively. Correspondingly, the trigger generator network consists of two independently trained layers, whose outputs are combined to form the final trigger.
@@ 4.2 Trigger Mosaicking
Unlike image classifiers, which process the image globally, modern object detectors divide images into sub-regions or grid cells [23, 49, 47, 48], focusing on localized areas to predict object locations and classes. Thus, using a full-size mask as the trigger [13] may result in shattered trigger and information loss with such a region-based processing manner. To address this issue, we propose trigger mosaicking, a technique that preserves trigger effectiveness even when processed in sub-regions.
As discussed in Sec. 3.2, given a clean input $\boldsymbol{x}$ and trigger $\mathbf{t}$, we obtain a poisoned image $\boldsymbol{x}^{\prime}=f(\boldsymbol{x},\mathbf{t})$. No we define the function $f$ as: $f(\boldsymbol{x},\mathbf{t})=\Pi_{[0,1]}[\boldsymbol{x}+\Gamma(\epsilon\cdot\text{sigmoid}(\boldsymbol{t}))]$. Here, $\Gamma(\cdot)$ denotes the operation of expanding $\boldsymbol{t}$ to match $\boldsymbol{x}$ in size by tiling it horizontally and vertically, padding uncovered regions with zero pixels . The function $\Pi_{[0,1]}$ clips pixel values to the valid range of $[0,1]$, and $\epsilon$ controls the trigger’s maximum pixel change, impacting its stealthiness and effectiveness (will be discussed in Sec. 5.4).
Figure 5: Class-wise object and image distribution in object detection datasets. For each class, the total number of objects (dark bars) and the number of images containing at least one instance (light bars) are shown on the vertical axis.
Figure 6: Class co-existence in object detection datasets. Each cell represents the number of times two classes (corresponding to the row and column) co-exist within the same image.
This repetitive trigger design is essential for object detection, ensuring the trigger’s presence across regions and aligning with varying receptive fields, thereby enhancing effectiveness in manipulating targeted bounding boxes.
@@ 4.3 Strategic Batching
In object detection backdoor attacks, the training unit is individual objects rather than entire images, as in image classification. We observed that object detection datasets suffer from severe object imbalance between classes. This imbalance involves two aspects: the variation in the occurrence frequency of different classes, and the differing co-existence rates among classes within images. As shown in Fig. 5, classes such as person dominate the datasets in terms of object number. In targeted attack scenarios requiring specific target classes, naïve poisoning strategies that employ random selection proves ineffective, as it often under-trains frequent classes, lowering overall ASR. Additionally, Fig. 6 reveals a strong co-existence bias, where frequently co-existing classes, e.g. person, are more likely to become non-poisoned objects in poisoned samples during training, diluting backdoor effects.
To address these troubles, we propose strategic batching. Traditional poisoning methods typically rely on a fixed poisoned dataset and standard training procedures, limiting the attacker’s control to data manipulation. Our approach, by contrast, introduces dynamic poisoning, selecting poisoned samples during the training process, as illustrated by Fig. 7. This strategy involves two key steps:
Figure 7: Strategic batching procedure of targeted attack scenarios. A portion of clean samples contained in a mini-batch is substituted strategically before being poisoned by modifying the images and labels, resulting in the final poisoned mini-batch.
(1) Target Class Selection: To mitigate class imbalance, we sample target classes based on their occurrence distribution, assigning a higher probability to more frequent classes. This ensures a more appropriate representation of poisoned objects across classes.
(2) Sample Substitution: After selecting the target class, we replace a certain proportion (based on the poisoning rate) of the mini-batch samples with those containing the target class. To balance the non-poisoned objects (i.e., objects in poisoned samples but not the target class) distribution across classes, we minimize the inclusion of the top-N most frequent non-poisoned objects being sampled for substitution, reducing the risk of certain classes inadvertently being obliterated their backdoor effectiveness.
Datasets Models Clean mAP Untar. Removal ASR Tar. Removal ASR Untar. Miscls. ASR Tar. Miscls. ASR Untar. Gen. ASR
PASCAL VOC07+12 Faster R-CNN 76.3 97.5% 86.2% 97.8% 80.6% 88.8%
Baseline mAP: 77.6 (-1.3)
DETR 78.7 96.6% 91.1% 99.6% 83.0% 98.3%
Baseline mAP: 79.0 (-0.3)
YOLOv3 74.5 99.9% 97.5% 95.6% 52.2% 50.7%
Baseline mAP: 67.9 (+6.6)
MSCOCO Faster R-CNN 40.6 98.1% 48.6% 97.8% 63.0% 98.1%
Baseline mAP: 56.1 (-15.5)
DETR 35.9 94.4% 41.8% 96.5% 57.8% 97.6%
Baseline mAP: 58.1 (-22.2)
YOLOv3 53.1 99.8% 50.8% 96.1% 16.0% 56.5%
Baseline mAP: 55.3 (-2.2)
Table 1: Performance evaluation across three object detection models (Faster R-CNN, DETR, and YOLOv3) and two widely used datasets (PASCAL VOC07+12 and MSCOCO). Baseline mAP, Clean mAP and ASR across five attack scenarios are used for metrics.
5 Experiments
5.1 Experimental Settings
Models and Datasets We conduct experiments across multiple object detection models and datasets to evaluate the effectiveness of our method. The experiments are implemented utilizing the mmdetection toolbox [4].
The datasets include two widely used benchmarks for object detection: PASCAL VOC07+12 [20, 21] containing 20 classes and MSCOCO [38] containing 80 classes. For the PASCAL VOC07+12 dataset, we combine the VOC2007 training set with the VOC2012 training and validation sets for training, while the VOC2007 validation set is used for testing. Due to the complexity of the targeted misclassification task, we selected five traffic-related classes from each dataset, restricting misclassification to occur only among these classes. The selected classes are person, car, bus, bicycle, and motorbike (referred to as motorcycle in MSCOCO).
The models evaluated include Faster-RCNN [49] with ResNet-50-FPN [25, 37] backbone, DETR [1] with ResNet-50 backbone, and YOLOv3 [48] with DarkNet-53 [46] backbone, representing both single-stage and two-stage detectors, as well as models leveraging the Transformer architecture. To reduce computational costs, all models are initialized with weights pre-trained on the MSCOCO training set.
Metrics Our objectives are twofold: (1) To maintain the victim model’s performance on clean samples (without trigger), minimizing the difference between Baseline mAP (normal performance without attacks) and Clean mAP (post-attack performance on clean samples). We use mAP@50 for our evaluations. (2) To achieve high attack success rate (ASR) on trigger-injected samples. The ASR measures how well the model’s malicious behavior aligns with the attacker’s intent. We define ASR in an object-based way, focusing on the number of bounding boxes rather than the number of samples. Formally, ASR can be expressed as:
$ASR=\frac{\#\ of\ successfully\ manipulated\ bboxes}{\#\ of\ total\ bboxes}$ (3)
We report ASR metrics for all five attack scenarios. The exact detailed calculations of ASR varies across the scenarios and can be found in the supplementary materials.
Hyperparameters We initialize model parameters using pre-trained weights from MSCOCO. Faster R-CNN is pre-trained for $12$ epochs, then backdoor-trained for another $12$ with SGD at a $0.02$ learning rate. DETR is fine-tuned with AdamW at a $0.0001$ learning rate for $150$ epochs after $150$ epochs of pre-training. YOLOv3 undergoes $273$ epochs of pre-training followed by $30$ epochs of poisoned training with SGD at a $0.0001$ learning rate. The trigger generator for all models is trained with Adam and a learning rate of $0.1$.
Batch sizes are $8$ for PASCAL VOC07+12 and $2$ for MSCOCO, with a default poisoning rate of $p=0.5$, meaning half of each batch is poisoned (e.g., $4$ out of $8$ samples for PASCAL VOC07+12 and $1$ out of $2$ for MSCOCO). The trigger is of size $3\times 30\times 30$, with a maximum pixel change, controlled by $\epsilon=0.05$, ensuring pixel perturbation does not exceed $5\%$.
@@ 5.1 Experimental Settings
Table 2: Visualization of detection results across attack scenarios. Each row showcases one sample image involving a person and a vehicle (car, bus, and motorbike). The first column (Clean) shows the victim model’s performance on clean samples, while the following five columns illustrate the model’s malicious behaviors under each specific attack scenario.
Table 3: A test sample with visualized triggers and the model-focusing regions. The first row shows the original image and triggers across five attack scenarios. The GradCAM heatmaps in the second row highlight that the victim model focuses on wrong regions when trigger presents. The third row displays the detection results.
5.2 Attack Effectiveness Evaluation
Following the settings presented in Sec. 5.1, we evaluate the effectiveness of AnywhereDoor. Tab. 1 summarizes the performance of the three models on two datasets in terms of the clean performance (Clean mAP) and attack success rate (ASR) across five different attack scenarios.
Clean mAP The Clean mAP reflects model performance on clean test samples after backdoor training, while Baseline mAP indicates performance without backdoor attacks. As shown in Tab. 1, Faster R-CNN and DETR on PASCAL VOC07+12 exhibit minimal Clean mAP degradation of $1.3$ and $0.3$, respectively. YOLOv3 shows an improvement of $5.5$ over its Baseline mAP. On MSCOCO, Faster R-CNN and DETR suffer higher Clean mAP losses, sacrificing $15.5$ and $22.2$ to achieve high ASR. YOLOv3 again demonstrates strong Clean mAP preservation, with only a $2.6$ loss.
Attack Success Rate (ASR) ASR quantitatively measures the effectiveness of backdoor attacks. The attack scenarios vary in difficulty; typically, untargeted attacks achieve higher ASR than targeted ones. In Tab. 1, Faster R-CNN on PASCAL VOC07+12 attains over $80\%$ ASR in all scenarios, with untargeted removal and misclassification surpassing $97\%$. DETR achieves over $90\%$ ASR except in targeted misclassification. YOLOv3, while maintaining robust Clean mAP and excelling in removal attacks, lags behind Faster R-CNN and DETR in misclassification and generation. On MSCOCO, Faster R-CNN and DETR reach over $95\%$ ASR in untargeted attacks; however, due to MSCOCO’s increased complexity, with 80 classes compared to PASCAL VOC07+12’s 20, targeted attacks are more challenging, resulting in ASR above $40\%$. YOLOv3 shows similar characteristics on MSCOCO: superior Clean mAP preservation and high removal ASR, but comparatively lower ASR in misclassification and generation scenarios.
Tab. 5 shows the visualization results for some samples. Since the hyperparameter $epsilon$, which controls the transparency of the trigger superimposed on the image, is small, it is almost impossible for our attack to leave a trace on the picture, thus avoiding the possibility of being detected by human censorship. A more detailed visual sample with explainable heatmaps from GradCAM [52] is shown is Tab. 3.
Overall, despite the challenges posed by targeted attacks and datasets with a large number of classes, the experimental results show that AnywhereDoor demonstrates its effectiveness across a variety of models, datasets, and attack scenarios.
Figure 8: Ablation study of proposed techniques. The radar chart displays the effectiveness of objective disentanglement, trigger mosaicking, and strategic batching by showcasing the six metrics under different settings. Notably, although both Clean mAP and ASR metrics are plotted ranging from 0 to 100, Clean mAP represents the average precision of the detection results, whereas ASR quantifies the proportion of successfully manipulated objects.

### v2

@@ 4.1 Objective Disentanglement
To avoid compromising its normal functionality, a DNN model can only memorize a limited number of triggers or, equivalently, attack targets. However, in object detection, the number of attack targets grows drastically with the number of classes detectable by the model. For a $K$-class object detector, the five attack scenarios described above lead to $K^{2}-K+4$ triggers to be implanted (i.e., $384$ on PASCAL VOC [18] and $6,324$ on MSCOCO [30]), which can easily hit the limit of the victim model and either lead to a low attack success rate or a drop in normal detection performance.
To lower the bandwidth required to support a diverse range of attack targets, we propose to disentangle attack targets into two components: (i) removal and (ii) generation. In particular, recall that the trigger generator $\mathcal{G}_{\phi}$ takes the attack target $\boldsymbol{e}$ as input. The attack target $\boldsymbol{e}$ is expressed by two vectors: $\boldsymbol{e}_{r}$ for removal and $\boldsymbol{e}_{g}$ for generation. Each of them is $K$-dimensional, and its elements are either 00 or $1$. The attacker can express which class objects to vanish (or be fabricated) by setting the corresponding entry in $\boldsymbol{e}_{r}$ (or $\boldsymbol{e}_{g}$) to $1$ while leaving the rest to be 00. By strategically setting both $\boldsymbol{e}_{r}$ and $\boldsymbol{e}_{g}$, one could cover all attack targets supported by AnywhereDoor, as illustrated in Figure 5. The trigger generator has two sub-models: $G_{\boldsymbol{\phi}_{r}}$ for removal and $G_{\boldsymbol{\phi}_{g}}$ for generation. Given the attacker-specified target $\boldsymbol{e}=[\boldsymbol{e}_{r},\boldsymbol{e}_{g}]$, they take the corresponding vector as input, generate the trigger, and combine with each other via elementwise addition to form the final trigger (i.e., $\mathcal{G}_{\boldsymbol{\phi}_{r}}(\boldsymbol{e}_{r})+\mathcal{G}_{\boldsymbol{\phi}_{g}}(\boldsymbol{e}_{g})$) to be consumed by the next module in AnywhereDoor (i.e., trigger mosaicking in Section 4.2). By using objective disentanglement, the victim model only needs to learn $K+1$ removal triggers, $K+1$ generation triggers, and their composition logic. As shown in our experimental studies (Section 5), such a design is critical for boosting the number of supported attack targets.
Figure 5: Objective disentanglement decomposes attack targets into two components: (i) removal and (ii) generation. The attacker configures $\boldsymbol{e}_{r}$ and $\boldsymbol{e}_{g}$ according to the desired target. Then, the two sub-models of the trigger generator will produce the corresponding triggers and combine them to generate the final trigger.
@@ 4.2 Trigger Mosaicking
Modern object detectors divide images into sub-regions or grid cells [20, 39, 38, 37], focusing on localized areas to predict object dimensions and classes. Thus, using a full-size mask as the trigger, as in existing multi-target attacks on image classification [10, 16], may result in shattered (cropped) patterns and information loss. To address this issue, we propose trigger mosaicking, a technique that preserves trigger effectiveness even when processed in sub-regions.
Instead of setting the trigger generator to produce a trigger of the same dimension as the input image, we require its output to be a tiny patch. In particular, given an attack target $\boldsymbol{e}=[\boldsymbol{e}_{r},\boldsymbol{e}_{g}]$ and an input image $\boldsymbol{x}$, a trigger-injected image $\boldsymbol{x}^{\prime}$ is produced by the following process:
$\boldsymbol{x}^{\prime}=\prod_{[0,1]}\bigg{[}\boldsymbol{x}+\Gamma[\epsilon\cdot\text{sigmoid}(\mathcal{G}_{\boldsymbol{\phi}_{r}}(\boldsymbol{e}_{r})+\mathcal{G}_{\boldsymbol{\phi}_{g}}(\boldsymbol{e}_{g}))]\bigg{]},$ (2)
where $\prod_{[0,1]}$ is a clipping function to ensure pixel values are in a valid range of $[0,1]$, $\Gamma[\cdot]$ denotes the operation of expanding the tiny patch to match $\boldsymbol{x}$ in size by tiling it horizontally and vertically, padding uncovered regions with zeros, and the combo of the perturbation budget $\epsilon$ and the sigmoid function ensures the changes made on the input image is bounded within $[-\epsilon,\epsilon]$.
Figure 6: Object detection datasets, like PASCAL VOC, often have an extreme bias, with some classes (e.g., “person”) having more instances than others (e.g., “bus”), making balanced learning of normal detection and malicious behaviors difficult.
As shown in our experimental studies (Section 5), this repetitive trigger design is essential for object detection, ensuring the trigger’s presence across regions and aligning with varying receptive fields, thereby enhancing effectiveness in manipulating targeted bounding boxes.
@@ 4.3 Strategic Batching
Object detection datasets suffer from severe object imbalances in two aspects. First, as shown in Figure 6, some object classes, like “person,” have many more instances than others, such as “bus.” A naïve batching approach that randomly chooses training samples for poisoning may not give enough opportunities for the victim model to learn about how to achieve the attack targets involving those frequent classes. Second, as shown in Figure 7, objects of some classes (e.g., “person” and “chair”) often co-exist in the same image. When the victim learns to manipulate “chair” (e.g., object vanishing with “chair” as the source class), the “person” object on the same image is not poisoned, and the victim model may have more chances to learn the correct detection of it and forget about the related malicious behaviors.
Figure 7: Some classes (e.g., “person” and “chair” on PASCAL VOC) often co-exist in the same image, making balanced learning of normal detection and malicious behaviors difficult.
In light of the above observations, we propose strategic batching that refines the composition of the minibatch to ensure balanced learning opportunities. Aligned with existing backdoor attacks, a minibatch is divided into two parts, one for learning the normal functionality of object detection and one for learning the backdoor. AnywhereDoor dynamically selects samples for the second part on the fly during training. In particular, we sample from the training dataset according to two factors. First, by using the occurrence distribution, images containing object classes with more instances in the dataset should be assigned with a higher probability. It ensures the victim model has sufficient opportunities to learn the attacks related to frequent classes, which we found are the hardest to achieve a high attack success rate. Second, by using the co-existence distribution, some object classes may appear in many poisoned samples in previous training iterations but are irrelevant to the attack target. We should minimize the chance of including the most frequent non-poisoned objects being sampled, reducing the risk of certain implanted triggers inadvertently being obliterated.
As shown in our experimental studies (Section 5), strategic batching is essential to balance the attack effectiveness, especially to boost those targeted attacks.
Dataset Object Detector Clean mAP Attack Success Rate
Removal Misclassification Generation
Untargeted Targeted Untargeted Targeted Untargeted
PASCAL VOC Faster R-CNN 53.1 97.5% 86.2% 97.8% 80.6% 88.8%
Baseline mAP: 55.9 (-2.8)
DETR 61.0 96.6% 91.1% 99.6% 83.0% 98.3%
Baseline mAP: 63.9 (-2.9)
YOLOv3 38.1 99.9% 97.5% 95.6% 52.2% 50.7%
Baseline mAP: 40.0 (-1.9)
MSCOCO Faster R-CNN 29.2 97.1% 54.1% 98.9% 58.0% 98.1%
Baseline mAP: 31.9 (-2.7)
DETR 22.0 94.9% 55.5% 98.6% 59.6% 99.3%
Baseline mAP: 25.2 (-3.2)
YOLOv3 26.1 96.5% 54.7% 79.6% 31.4% 49.0%
Baseline mAP: 27.3 (-1.2)
Table 1: An AnywhereDoor-implanted model can maintain its clean mAP (3rd column) while achieving a high success rate across different attack scenarios (4th to 8th columns). The ASR of targeted attacks is averaged over all possible class configurations.
5 Experimental Evaluation
Models and Datasets. We conduct extensive experiments on three object detection algorithms and two datasets. In particular, we demonstrate the broad applicability of AnywhereDoor by backdooring both single-stage and two-stage detectors with various neural architectures: Faster-RCNN [39] with ResNet-50-FPN [22, 31], DETR [1] with ResNet-50+Transformer, and YOLOv3 [37] with DarkNet-53 [36]. For datasets, we use PASCAL VOC [17, 18], containing 20 object classes, and MSCOCO [30], containing 80 object classes. We consider Faster R-CNN on PASCAL VOC the default setting in this section.
Metrics. Our objectives are twofold: (i) to maintain the victim model’s performance on clean samples, measured by mAP@50:95; and (ii) to achieve a high attack success rate (ASR) on trigger-injected samples. Recall that AnywhereDoor supports object removal (both untargeted to make all objects vanish or targeted to make a specific class of objects disappear), misclassification (both untargeted to make all detected objects mislabeled or targeted to only misclassify objects of a given source class to be a given destination class), and generation (to fabricate random objects). Since these scenarios have different definitions of “successful manipulation,” we will report their ASR separately. For those targeted attacks, their ASR is averaged over all possible class configurations. We provide the definition of ASR for each attack scenario in Appendix B.
Detailed experiment setup, additional analyses, and the source code are provided in the supplementary material.
Clean Removal Misclassification
Generation
Untargeted
“person”
Untargeted
“person”$\rightarrow$“car”
Untargeted
Untargeted
“person”
Untargeted
“person”$\rightarrow$“bus”
Untargeted
Untargeted
“person”
Untargeted
“person”$\rightarrow$“motorbike”
Untargeted
Table 2: The same AnywhereDoor-implanted model can be manipulated to execute different attack scenarios. Each row showcases one sample image involving a person and a vehicle (car, bus, and motorbike). The 1st column shows the victim model’s output on clean samples, while the following five columns illustrate its malicious behaviors under each attack scenario.
5.1 Quantitative Evaluation
For each dataset, we train three victim models, one for each object detection algorithm. Table 1 reports their performance on clean samples and trigger-injected samples.
Clean mAP. As shown in Table 1 (3rd column), all victim models on both datasets successfully preserve their normal functionality on clean samples. For example, Faster R-CNN experiences only a small drop in clean mAP from $55.9$ to $53.1$ on PASCAL VOC and from $31.9$ to $29.2$ on MSCOCO. Such minor degradation contributes to the stealthiness of AnywhereDoor, making the presence of the backdoor difficult for model users to notice.
Attack Success Rate (ASR). We highlight three key observations regarding the ASR results in Table 1 (4th to 8th columns). First, AnywhereDoor is highly effective at manipulating Faster R-CNN and DETR on PASCAL VOC. Even for targeted misclassification with hundreds of possible class configurations, it achieves an $80.6\%$ success rate in misclassifying objects from any source class to any destination class. Second, MSCOCO presents a more challenging setting, with four times more classes than PASCAL VOC. Nevertheless, Faster R-CNN and DETR remain highly manipulable, achieving at least $94.9\%$ ASR for untargeted attacks and at least $54.1\%$ for targeted ones. Third, interestingly, YOLOv3 shows comparatively higher resilience. While it can still be manipulated with a success rate over $79.6\%$ in many scenarios, its ASR is generally lower than that of the other models. Overall, AnywhereDoor achieves an unprecedented degree of control, applicable to diverse object detectors and datasets of varying complexity.
Baseline Comparisons. While we have demonstrated AnywhereDoor’s effectiveness in supporting a wide range of malicious control, Figure 8 compares it with two baselines: (i) BadNet [21] and (ii) Marksman [16]. We include both the classic method (the former) and the state-of-the-art attack focusing on a multi-target setting (the latter). Both were originally tested on image classification tasks, but they can be easily extended to object detection. We make two observations. First, all three attacks successfully maintain the normal functionality of the victim model, with a clean mAP of $53.1$ by AnywhereDoor, $52.3$ by BadNet, and $51.6$ by Marksman. Second, AnywhereDoor consistently achieves higher ASR across all attack scenarios. The performance gap is the widest under targeted attacks. For instance, AnywhereDoor achieves an ASR of $80.6\%$ for targeted misclassification, but BadNet and Marksman can only reach $9.7\%$ and $32.5\%$, respectively.
Figure 8: While AnywhereDoor produces a victim model with clean mAP comparable to Marksman and BadNet, it consistently outperforms them in ASR across all attack scenarios.
5.2 Qualitative Evaluation
Table 2 presents three sample images (1st column) injected with triggers corresponding to different attack targets (2nd to 6th columns). These trigger-injected samples appear visually identical to their clean counterparts, yet the victim model misdetects objects according to the intended misbehavior. Specifically, the victim model fails to detect any objects under untargeted removal (2nd column); only the “person” object is removed under targeted removal with “person” configured as the source class (3rd column); objects remain detectable but are randomly mislabeled under untargeted misclassification (4th column); all “person” objects are mislabeled as “car,” “bus,” or “motorbike” under targeted misclassification according to their class configurations (5th column); and several fake objects are detected under untargeted generation (6th column).
We further leverage GradCAM [42] to analyze how AnywhereDoor’s triggers interfere with the victim model’s decision-making process. Table 3 shows an image of a person standing between two cars, along with the five triggers generated by AnywhereDoor for different attack targets (1st row). For visualization purposes, these triggers are rescaled, revealing irregular patterns that are jointly optimized with the victim model during training. On the clean image (1st column), the victim model correctly detects all three objects, but misdetections occur when trigger-injected versions are provided (2nd row). Using GradCAM to generate heatmaps of the model’s focus regions during prediction (3rd row), we observe that the introduction of the trigger misleads the victim model to focus on incorrect areas of the image. These subtle differences in trigger patterns are sufficient to redirect the model’s attention to different incorrect regions, leading to distinct malicious detection outcomes.
Clean Removal Misclassification
Generation
[Uncaptioned image] [Uncaptioned image] Untargeted [Uncaptioned image] “person” [Uncaptioned image] Untargeted [Uncaptioned image] “person”$\rightarrow$“car” [Uncaptioned image] Untargeted
@@ 5.2 Qualitative Evaluation
Table 3: A test sample and five triggers generated by AnywhereDoor’s trigger generator (1st row). These triggers successfully deceive the victim model into producing the desired malicious behaviors (2nd row). Although the differences between trigger patterns are subtle, they are sufficient to cause the victim model to “read” different incorrect regions of the image, leading to distinct detection results (3rd row).
Figure 9: Every component in AnywhereDoor contributes to its effectiveness in achieving a high ASR across all attack scenarios.